The ages of employees at a particular company have a mean of 39. The distribution of sample means for samples of size 200 is normal with a mean of 39 and a standard deviation of 0.79. Suppose you take a sample of size 200 employees from the company and find that their mean age is 41.3. How many standard deviations is the sample mean above the mean of the sampling distribution?
the answer is 2.9 please show work

Answer by stanbon(75887)

The ages of employees at a particular company have a mean of 39.
The distribution of sample means for samples of size 200 is normal with a mean of 39 and a standard deviation of 0.79.
---------------------
Suppose you take a sample of size 200 employees from the company and find that their mean age is 41.3.
---------------------
How many standard deviations is the sample mean above the mean of the sampling distribution?2
z(41.3) = (41.3-39)/0.79 = 2.3/0.79 = 2.91
-----------------
Cheers,
Stan H.
----------------
the answer is 2.9 please show work